Status Report

Hubble Space Telescope Daily Report #2982 – 23 Oct 2001

By SpaceRef Editor
October 23, 2001
Filed under , ,

HUBBLE SPACE TELESCOPE

DAILY REPORT #2982

PERIOD COVERED: 0000Z (UTC) 10/22/01 – 0000Z (UTC) 10/23/01

Daily Status Report as of 296/0000Z

1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED:

1.1 Completed Two Sets of STIS/CCD 8901 (Dark Monitor-Part 1)

The Space Telescope Imaging Spectrograph (CCD) was used to monitor
the darks. There was no anomalous activity.

1.2 Completed Five Sets of WF/PC-2 8936 (Cycle 10 Supplemental Darks Pt1/3)

The WF/PC-2 was used to perform a dark calibration program that
obtains three dark frames every day to provide data for monitoring and
characterizing the evolution of hot pixels. The proposal completed with no
reported problems.

1.3 Completed STIS/CCD 9128 (A Snapshot Survey of Absorption Lines from
High Velocity Clouds in the Milky Way)

The Space Telescope Imaging Spectrograph (CCD) was used to search
for Mg II and Mg I absorption from Galactic High Velocity Clouds {HVCs}
along the sightlines of 51 of the brightest QSOs, BL Lacs and Seyfert 2
galaxies in the sky. The observations completed with no reported problems.

1.4 Completed Four Sets of WF/PC-2 8935 (Cycle 10 Standard Darks)

The WF/PC-2 was used to obtain dark frames every week in order to
provide data for the ongoing calibration of the CCD dark current rate, and
to monitor and characterize the evolution of hot pixels. Over an extended
period these data will also provide a monitor of radiation damage to the
CCDs. The proposal completed with no reported problems.

1.5 Completed Three Sets of STIS/CCD/MA1/MA2 9116 (Understanding
High-Redshift and Starburst Galaxies: A UV Spectroscopic Survey of B- Stars
in the SMC)

The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was
used to construct a high quality spectral library at low metallicity in
order to synthesize the UV spectra of high redshift star-forming galaxies
and nearby starbursts. No problems were reported.

1.6 Completed Three Sets of WF/PC-2 8059 (POMS Test Proposal: Targeted
Parallel Archive Proposal)

The WF/PC-2 was used to observe the parallel opportunities
available in the neighborhood of bright galaxies are treated in a slightly
different way from the normal pure parallels. Local Group galaxies offer
the opportunity for a closer look at young stellar
populations. Narrow-band images in F656N can be used both to identify
young stars via their emission lines, and to map the gas distribution in
star-forming regions. The observations completed nominally.

1.7 Completed WF/PC-2 8939 (Cycle 10 Internal Monitor)

The WF/PC-2 was used to calibrate the internal monitor, to be run
weekly to monitor the health of the cameras. No problems were encountered.

1.8 Completed STIS/CCD 8903 (Bias Monitor – Part 1)

The Space Telescope Imaging Spectrograph (CCD) was used to monitor
the bias in the 1×1, 1×2, 2×1, and 2×2 bin settings at gain=1, and 1×1 at
gain = 4 to build up high-S/N superbiases and track the evolution of hot
columns. There were no problems.

1.9 Completed STIS/CCD 8262 (Nuclear Kinematics of the Dense Globular
Cluster M-15)

The Space Telescope Imaging Spectrograph (CCD) was used to make
observations of star NGC7078-CE as part of the study of the nuclear
kinematics of the dense globular cluster M-15. The high central density of
M-15 makes it unique for probing the structure and evolution of globular
clusters. The observations were completed as planned, and no anomalies
were reported.

1.10 Completed Three Sets of WF/PC-2 9244 (POMS Test Proposal: WFII
Parallel Archive Proposal Continuation)

The WF/PC-2 was used to perform a generic target version of the
Archival Pure Parallel program. The program will be used to take parallel
images of random areas of the sky, following the recommendations of the
Parallels Working Group. The observations completed with no anomalous
activity.

1.11 Completed STIS/CCD 9066 (Closing in on the Hydrogen Reionization
Edge of the Universe)

The Space Telescope Imaging Spectrograph (CCD) was used in parallel
constrain the Hydrogen reionization edge in emission that marks the
transition from a neutral to a fully ionized IGM at a predicted
redshifts. The proposal completed uneventfully.

1.12 Completed WF/PC-2 9085 (Measuring The Masses Of High-Z Quasar Host
Galaxies)

The WF/PC-2 was used to investigate the close link between
black-hole mass and spheroid mass in both quiescent and active galaxies in
the local universe. Consequently black-hole and spheroid formation/growth
are now viewed as intimately related processes, and establishing the mass
of quasar host galaxies as a function of redshift is now seen as a key
measurement in observational cosmology. From the Cycle 7 NICMOS program,
the best estimate to date of the mass evolution of the hosts of both
radio-quiet and radio-loud quasars out to z ~eq 2 has been derived. Under
the assumption of passive stellar-population evolution our results are
consistent with the black-hole/spheroid population being unchanged out to z
~eq 2. However, the crucial assumption of passive evolution needs to be
tested. All observations competed nominally.

1.13 Completed STIS/CCD 9285 (POMS Test Proposal: STIS Non-scripted
Parallel Proposal Continuation III)

The Space Telescope Imaging Spectrograph (CCD) was used to make low
galactic latitude, non-scripted parallel observations as part of a POMS
test proposal. The observations were completed as planned, and no
anomalies were reported.

1.14 Completed WF/PC-2 9160 (Disks and Envelopes of Nearby Nebulous
Young Stellar Objects: A Snapshot Survey)

The WF/PC-2 was used to perform a snapshot survey of nearby
nebulous young stellar objects to study the detailed morphology of the
their disks and envelopes and probe the effect of inclination on the
infrared spectral energy distribution of disk/envelope systems. The
proposal completed as planned.

2.0 FLIGHT OPERATIONS SUMMARY:

2.1 Guide Star Acquisitions:

Scheduled Acquisitions: 6

Successful: 6

Scheduled Re-acquisitions: 7

Successful: 7

2.2 FHST Updates:

Scheduled: 9

Successful: 9

2.3 Operations Notes:

GSFC Internal Simulation #14 (Planning/Contingency EVA) ran
yesterday from 8:00 a.m. until approximately 3:30 p.m.. It will continue,
supported by many operations personnel, today from 8:00 a.m. until 5:00 p.m.

Using ROP SR-1A, the SSR EDAC error counter was cleared three times.

SSA transmitter #1 was turned on and off for the interval 295/2304Z
to 295/2317Z as directed by ROP IC-2.

The ephemeris table was uplinked at 296/082714Z per ROP DF-07A.

3.0 SIGNIFICANT FORTHCOMING EVENTS:

Continuation of normal science observations and calibrations.

SpaceRef staff editor.